# AI-Driven Fuzzing for Vulnerability Assessment of 5G Traffic Steering Algorithms

### v1

@@ Abstract
Traffic Steering (TS) dynamically allocates user traffic across cells to enhance Quality of Experience (QoE), load balance, and spectrum efficiency in 5G networks. However, TS algorithms remain vulnerable to adversarial conditions such as interference spikes, handover storms, and localized outages. To address this, an AI-driven fuzz testing framework based on the Non-Dominated Sorting Genetic Algorithm II (NSGA-II) is proposed to systematically expose hidden vulnerabilities. Using NVIDIA Sionna, five TS algorithms are evaluated across six scenarios. Results show that AI-driven fuzzing detects 34.3% more total vulnerabilities and 5.8% more critical failures than traditional testing, achieving superior diversity and edge-case discovery. The observed variance in critical failure detection underscores the stochastic nature of rare vulnerabilities. These findings demonstrate that AI-driven fuzzing offers an effective and scalable validation approach for improving TS algorithm robustness and ensuring resilient 6G-ready networks.
@@ I Introduction
To address this gap, this study proposes an artificial intelligence (AI)-driven fuzzing framework based on NSGA-II to generate adversarial network states that expose rare TS failures affecting QoE, fairness, and stability. The approach outperforms traditional testing by 34.3% in overall vulnerability detection, with p << 0.00001 indicating a highly significant improvement unlikely due to chance. It further provides architecture-level insights, showing that rule-based TS policies are more prone to failures under adversarial conditions than adaptive learning-based agents. These results establish AI-driven fuzzing as a statistically validated and architecture-aware method for enhancing the robustness of 5G TS systems.
@@ II Background and Related Work
A vulnerability denotes network states where performance degrades beyond acceptable limits (per 3GPP TS 22.261), affecting stability, QoE, or fairness. Current testing methodologies for TS algorithms fall into three categories: Conventional testing, which uses predefined scenarios but misses edge cases, formal verification, which provides theoretical guarantees but becomes computationally intractable for realistic 5G models with hundreds of state variables [1], and fuzzing techniques, which have shown success for protocol testing [10, 12, 11] and which focus on implementation errors rather than on algorithmic vulnerabilities specific to TS logic. The gap between optimization-focused TS research and security-oriented fuzzing presents an opportunity. Multi-objective genetic algorithms have explored network tradeoffs [2], but have not be applied for vulnerability discovery. This study bridges this gap by adapting NSGA-II for TS vulnerability assessment, introducing domain-specific fitness functions that simultaneously capture QoE degradation, fairness violations, and stability issues. Beyond academic relevance, the proposed framework also provides practical utility for telecom vendors’ back-office teams by enabling systematic identification of TS vulnerabilities during testing and validation.
III System Model, Problem Formulation, and Methodology
@@ III-B Problem Formulation
where $H_{\text{rate}}$ denotes the average handover rate per UE, $\text{Thr}_{5\%}$ is the 5th-percentile user throughput, and $J$ is Jain’s fairness index. The parameters $\tau_{\text{stab}}$, $\tau_{\text{QoE}}$, and $\tau_{\text{fair}}$ represent operator-defined thresholds.
@@
$f_{1}(z)$ $=\frac{1}{T}\sum_{t=1}^{T}(h_{t}-\bar{h})^{2}\quad$ $\text{(Instability)},$ (4)
@@
$f_{3}(z)$ $=1-\frac{(\sum_{t}u_{t})^{2}}{N\sum_{t}u_{t}^{2}}\quad$ $\text{(Unfairness)}.$ (6)
The formulation in (3)–(6) ensures simultaneous coverage of critical failure modes in 5G TS systems. Objective $f_{1}$ quantifies the temporal variance in the handovers (ping-pong instability), $f_{2}$ emphasizes performance degradation of the lower-tail users, and $f_{3}$ measures deviation from perfect resource fairness.
@@ IV-A NSGA-II-Based Vulnerability Discovery Engine
Each candidate configuration $z_{i}$ is evaluated through the objective vector $F(z_{i})=[f_{1},f_{2},f_{3}]$ defined in (3)–(6), capturing instability, QoE degradation, and unfairness. Non-dominated sorting and elitism ensure that superior solutions persist across generations, while the crowding distance metric maintains front diversity.
@@ V-A Simulation Setup
The evaluation encompasses six network scenarios and five TS algorithms using the Sionna simulator, with 300 runs per method (10 trials × 15 iterations). Table I details the simulation and algorithmic parameters.
@@ V-B Traffic Steering Algorithms Evaluated
The evaluation across six scenarios and five algorithms generated 14,223 vulnerability instances, revealing distinct patterns in how AI-Fuzzing and traditional testing discover network failures. Here, traditional testing refers to baseline scenario-based evaluations using fixed network configurations and deterministic test cases without AI-driven exploration. The results are analyzed across multiple dimensions: overall detection rate, severity distribution, algorithm-specific vulnerability, and convergence efficiency.
@@ V-C Vulnerability Discovery
Table II shows that AI-Fuzzing significantly outperforms traditional testing, detecting 34.3% more vulnerabilities ($t=8.671$, $p<0.00001$) and 5.8% more critical failures (200 vs 189). For total vulnerability counts, which exhibit a more symmetric distribution across runs, the parametric t-test is appropriate. However, for critical failures, which show extreme skewness (most runs detecting zero, with occasional high-count outliers), the non-parametric Mann-Whitney U test ($p=0.002$) provides more reliable evidence of superiority than the parametric t-test ($p=0.373$).
@@
Total Vulnerabilities 8,207 6,112
Vulnerabilities per Run 27.36 ± 8.59 20.37 ± 10.96
@@
Critical Failures per Run 0.67 ± 1.28 0.63 ± 1.49
Average Severity 3.24 2.28
Improvement Rate +34.3%
@@
Shannon Diversity 0.924 0.777
Cohen’s d (vulnerabilities) 0.708
@@
Figure 2: Vulnerability severity distribution across 300 runs (6 scenarios $\times$ 5 algorithms $\times$ 10 trials): AI-Fuzzing detects more total vulnerabilities (8,207 vs. 6,112, $p<0.00001$) and critical failures (200 vs. 189, $p=0.002$ Mann-Whitney U test) (a) . Normalized distribution shows superior diversity (Shannon: 0.924 vs. 0.777) with balanced severity coverage (b). High variance in critical failures (SD >> mean) reflects stochastic discovery of rare vulnerabilities.
@@ V-D Statistical Analysis and Confidence Intervals
AI-Fuzzing demonstrates significantly higher vulnerability detection than traditional testing, with a Shannon diversity index of 0.924 versus 0.777 and a strong effect size (Cohen’s d = 0.708). The 95% confidence intervals for critical failures per run are [0.52, 0.81] for AI-Fuzzing and [0.46, 0.80] for traditional testing, with high variance (SD = 1.28 and 1.49, respectively) reflecting the stochastic nature of critical failure discovery, where most test runs detect zero critical failures while occasional runs identify clusters of severe vulnerabilities. Due to strong right-skew (SD/mean ratios of 1.91 and 2.37, frequent zero counts violating normality assumptions), we prioritize the non-parametric Mann-Whitney U test ($p=0.002$) over the parametric t-test ($p=0.373$), as it is more robust to outliers and confirms AI-Fuzzing’s statistical advantage despite overlapping confidence intervals.
@@ V-G Scenario-Specific Performance Analysis
TABLE III: Vulnerability Detection Performance.
Traditional (Mean) AI-Fuzzing (Mean)
Scenario Value Improvement (%) Value p-value
@@
Stable Mobility 21.70 28.8 27.96 0.0000
Stable High Load 21.00 34.8 28.30 0.0000
Load Imbalance 20.70 45.4 30.10 0.0000
@@
Coverage Hole 21.18 30.1 27.56 0.0000
High Interference 21.52 42.3 30.62 0.0000
Congestion Crisis 16.14 21.4 19.60 0.0002
Average Improvement 33.8%
@@ V-H Practical Implications and Network Robustness
The results demonstrate clear advantages of AI-Fuzzing as the primary testing methodology for 5G network validation. AI-Fuzzing achieves statistically significant improvements across all metrics: 34.3% higher total vulnerability detection ($p<0.00001$, Cohen’s $d=0.708$) and 5.8% more critical failures ($p=0.002$ via Mann-Whitney U test). The more realistic critical failure threshold (>>3 UEs experiencing ping-pong, $\sim$7.5%) compared to initially conservative criteria enables sufficient statistical power, meaning a high probability of correctly detecting genuine performance differences.
@@ VI Conclusion
This paper establishes AI-driven fuzzing as superior to traditional testing for 5G TS vulnerability assessment. Using NSGA-II-based multi-objective optimization across six scenarios and five algorithms, AI-fuzzing achieves statistically significant improvements in both comprehensive coverage and critical edge-case detection (34.3% more vulnerabilities, $p<0.00001$), establishing it as the recommended primary validation approach for robust 5G/6G deployment. While simulation-based evaluation (7 cells, 40 UEs) enables controlled systematic assessment, future work will validate findings through Open RAN testbeds and develop GPU-accelerated adaptive fuzzing for production networks.
@@ Acknowledgment
This work was supported by the NTIA (Award No. 51-60-IF007) and NSF (Award No. CNS-2120442). The views expressed are those of the authors and do not necessarily represent those of the NTIA or NSF.

### v2

# Seyed Bagher Hashemi Natanzi1, Hossein Mohammadi2, Bo Tang1, Vuk Marojevic2

@@ Abstract
Traffic Steering (TS) dynamically allocates user traffic across cells to enhance Quality of Experience (QoE), load balance, and spectrum efficiency in 5G networks. However, TS algorithms remain vulnerable to adversarial conditions such as interference spikes, handover storms, and localized outages. To address this, an AI-driven fuzz testing framework based on the Non-Dominated Sorting Genetic Algorithm II (NSGA-II) is proposed to systematically expose hidden vulnerabilities. Using NVIDIA Sionna, five TS algorithms are evaluated across six scenarios. Results show that AI-driven fuzzing detects 34.2% more total vulnerabilities and 5.8% more critical failures than traditional testing, achieving superior diversity and edge-case discovery. The observed variance in critical failure detection underscores the stochastic nature of rare vulnerabilities. These findings demonstrate that AI-driven fuzzing offers an effective and scalable validation approach for improving TS algorithm robustness and ensuring resilient 6G-ready networks.
@@ I Introduction
To address this gap, this study proposes an artificial intelligence (AI)-driven fuzzing framework based on NSGA-II to generate adversarial network states that expose rare TS failures affecting QoE, fairness, and stability. The approach outperforms traditional testing by 34.2% in overall vulnerability detection, with p << 0.00001 indicating a highly significant improvement unlikely due to chance. It further provides architecture-level insights, showing that rule-based TS policies are more prone to failures under adversarial conditions than adaptive learning-based agents. These results establish AI-driven fuzzing as a statistically validated and architecture-aware method for enhancing the robustness of 5G TS systems.
@@ II Background and Related Work
A vulnerability denotes network states where performance degrades beyond acceptable limits (per 3GPP TS 22.261), affecting stability, QoE, or fairness. Current testing methodologies for TS algorithms fall into three categories: Conventional testing, which uses predefined scenarios but misses edge cases, formal verification, which provides theoretical guarantees but becomes computationally intractable for realistic 5G models with hundreds of state variables [1], and fuzzing techniques, which have shown success for protocol testing [9, 11, 10] and which focus on implementation errors rather than on algorithmic vulnerabilities specific to TS logic. The gap between optimization-focused TS research and security-oriented fuzzing presents an opportunity. Multi-objective genetic algorithms have explored network tradeoffs [2], but have not been applied for vulnerability discovery. This study bridges this gap by adapting NSGA-II for TS vulnerability assessment, introducing domain-specific fitness functions that simultaneously capture QoE degradation, fairness violations, and stability issues. Beyond academic relevance, the proposed framework also provides practical utility for telecom vendors’ back-office teams by enabling systematic identification of TS vulnerabilities during testing and validation.
III System Model and Problem Formulation
@@ III-B Problem Formulation
where $H_{\text{rate}}$ denotes the average handover rate per UE, $\text{Thr}_{5\%}$ is the 5th-percentile user throughput, and $J$ is Jain’s fairness index. The parameters $\tau_{\text{stab}}$, $\tau_{\text{QoE}}$, and $\tau_{\text{fair}}$ represent operator-defined thresholds aligned with 3GPP TS 22.261 service requirements to ensure operational relevance.
@@
$f_{1}(z)$ $=\frac{1}{N}\sum_{i=1}^{N}\mathcal{I}(c_{i}(z)\neq c_{i}^{\text{prev}})\quad$ $\text{(Instability)},$ (4)
@@
$f_{3}(z)$ $=1-\frac{\left(\sum_{i=1}^{N}T_{i}(z)\right)^{2}}{N\cdot\sum_{i=1}^{N}T_{i}(z)^{2}}\quad$ $\text{(Unfairness)}.$ (6)
The formulation in (3)–(6) ensures simultaneous coverage of 5G TS failure modes. Objective $f_{1}$ captures ping-pong instability, while $f_{2}$ targets QoE degradation using the base-2 logarithm throughput defined in (2). Objective $f_{3}$ measures unfairness via Jain’s Fairness Index. These objectives are evaluated against operator-defined thresholds $\tau_{\text{stab}}$, $\tau_{\text{QoE}}$, and $\tau_{\text{fair}}$, aligned with 3GPP TS 22.261 requirements, to robustly identify operational vulnerabilities.
@@ IV-A NSGA-II-Based Vulnerability Discovery Engine
Each candidate configuration $z_{i}$ is evaluated through the objective vector $F(z_{i})=[f_{1},f_{2},f_{3}]$ defined in (3)–(6), capturing handover instability, QoE degradation, and unfairness. Non-dominated sorting and elitism ensure that superior solutions persist across generations, while the crowding distance metric maintains front diversity.
@@ V-A Simulation Setup
The evaluation encompasses six scenarios and five TS algorithms using Sionna, totaling 300 runs per method (6 scenarios $\times$ 5 algorithms $\times$ 10 trials). To ensure a fair comparison, both AI-Fuzzing and traditional testing utilize a strictly matched computational budget and identical parameter ranges for network configuration sampling. Table I details the parameters.
@@ V-B Traffic Steering Algorithms Evaluated
The evaluation across six scenarios and five algorithms generated 14,319 vulnerability instances, revealing distinct patterns in how AI-Fuzzing and traditional testing discover network failures. Here, traditional testing refers to baseline scenario-based evaluations using fixed network configurations and deterministic test cases without AI-driven exploration. The results are analyzed across multiple dimensions: overall detection rate, severity distribution, algorithm-specific vulnerability, and convergence efficiency.
@@ V-C Vulnerability Discovery
Table II shows that AI-Fuzzing significantly outperforms traditional testing, detecting 34.2% more vulnerabilities ($t=8.671$, $p<0.00001$) and 5.8% more critical failures (200 vs 189). For total vulnerability counts, which exhibit a more symmetric distribution across runs, the parametric t-test is appropriate. However, for critical failures, which show extreme skewness (most runs detecting zero, with occasional high-count outliers), the non-parametric Mann-Whitney U test ($p=0.002$) provides more reliable evidence of superiority than the parametric t-test ($p=0.373$).
@@
Total Vulnerabilities 8,151 6,072
Vulnerabilities per Run 27.17 $\pm$ 8.24 20.24 $\pm$ 10.09
@@
Critical Failures per Run 0.67 $\pm$ 1.28 0.63 $\pm$ 1.49
Improvement Rate +34.2%
@@
Shannon Diversity 1.010 0.816
Cohen’s d (vulnerabilities) 0.752
@@
Figure 2: Vulnerability severity distribution across 300 runs (6 scenarios $\times$ 5 algorithms $\times$ 10 trials): AI-Fuzzing detects more total vulnerabilities (4,232 vs. 3,964, $p<0.00001$) and critical failures (200 vs. 189, $p=0.002$ Mann-Whitney U test) (a). Normalized distribution shows superior diversity (Shannon: 1.010 vs. 0.816) with balanced severity coverage (b). High variance in critical failures (SD >> mean) reflects stochastic discovery of rare vulnerabilities.
@@ V-D Statistical Analysis and Confidence Intervals
AI-Fuzzing demonstrates significantly higher vulnerability detection diversity than traditional testing, with a Shannon diversity index of 1.010 versus 0.816 and a medium-to-large effect size (Cohen’s $d=0.752$). The 95% confidence intervals for critical failures per run are [0.52, 0.81] for AI-Fuzzing and [0.46, 0.80] for traditional testing. The high variance (SD = 1.28 and 1.49, respectively) reflects the stochastic nature of critical failure discovery, where the majority of test runs detect zero critical failures while occasional runs identify clusters of severe vulnerabilities. Due to the strong right-skew (SD/mean ratios of 1.91 and 2.37) and frequent zero counts violating normality assumptions, we prioritize the non-parametric Mann-Whitney U test ($p=0.002$) over the parametric t-test ($p=0.373$). This robust analysis confirms AI-Fuzzing’s statistical superiority despite the overlapping confidence intervals.
@@ V-G Scenario-Specific Performance Analysis
TABLE III: Vulnerability Detection Performance across Scenarios.
Trad. AI-Fuzzing
Scenario Mean Imp. % Mean p-value
@@
Stable Mobility 21.70 28.8 27.96 0.0008
Stable High Load 21.00 34.8 28.30 0.0002
Load Imbalance 20.70 45.4 30.10 << 0.0001
@@
Coverage Hole 21.18 30.1 27.56 0.0004
High Interference 21.52 42.3 30.62 << 0.0001
Congestion Crisis 16.14 21.4 19.60 0.0463
Avg. Imp. 33.8%
@@ V-H Practical Implications and Network Robustness
The results demonstrate clear advantages of AI-Fuzzing as the primary testing methodology for 5G network validation. AI-Fuzzing achieves statistically significant improvements across all metrics: 34.2% higher total vulnerability detection ($p<0.00001$, Cohen’s $d=0.75$) and 5.8% more critical failures ($p=0.002$ via Mann-Whitney U test). The more realistic critical failure threshold (>>3 UEs experiencing ping-pong, $\sim$7.5%) compared to initially conservative criteria enables sufficient statistical power, meaning a high probability of correctly detecting genuine performance differences.
@@ VI Conclusion
This paper establishes AI-driven fuzzing as superior to traditional testing for 5G TS vulnerability assessment. Using NSGA-II-based multi-objective optimization across six scenarios and five algorithms, AI-fuzzing achieves statistically significant improvements in both comprehensive coverage and critical edge-case detection (34.2% more vulnerabilities, $p<0.00001$), establishing it as the recommended primary validation approach for robust 5G/6G deployment. While this study employs tabular Q-learning as a baseline representative, future work will extend the framework to Deep RL agents and RIC-based controllers within Open RAN testbeds to ensure scalability in production-grade 5G/6G networks.
@@ Acknowledgment
This work was supported in part by NTIA Awards 51-60-IF007 and 28-60-IF012 and NSF Award CNS-2120442. The views expressed are those of the authors and do not necessarily represent those of the NTIA or NSF.